# Efficient Spectral Element Method for the Euler Equations on Unbounded Domains in Multiple Dimensions

### v1

@@ 3.3 Constructing element matrices
$\psi_{l}(\mathbf{x})=h_{i}[\xi(\mathbf{x})]\otimes\overline{h}_{j}[\eta(\mathbf{x})],\hskip 5.0pt\hskip 3.0ptl=i+1+j(N_{LGL}),$ (29)
@@
$\mathbf{M}^{e}_{ij}$ $=\sum_{k=1}^{N+1}\sum_{m=1}^{N+1}\overline{\omega}(\xi_{k},\eta_{m})\psi_{i}(\xi_{k},\eta_{m})\psi_{j}(\xi_{k},\eta_{m})|\mathbf{J}(\xi_{k},\eta_{m})|,$ (31)
@@
$\omega_{\xi},\omega_{\eta}$ =$=\begin{cases}\omega,\omega&\textrm{ if $\Omega_{e}\in\Omega^{F}$}\\ \omega,\hat{\omega}&\textrm{ if $\Omega_{e}\in\Omega^{S}$}\end{cases}.$ (33)
@@ 3.4 Direct stiffness summation
To control unresolved grid-scale noise and aliasing, low-pass spectral filters are typically used with SEM [Sec. 18.3][13]. Spectral filters employ a three-step process to damp/remove unphysical high-frequency components: 1) the element-local nodal solution is transformed into modal space, 2) a low-pass filter is applied in modal space, 3) the filtered modal representation is inverse transformed to nodal space. The nodal to modal transform requires defining set of modal basis functions. As discussed in Section 3d of [14], the low-pass filter should not violate the continuity requirement of the SEM. One way to enforce this is by choosing modal functions such that most of the modes are zero at the element boundary. For a 1D element using LGL points, an appropriate choice is $\phi_{k}(\xi)=P_{k}(\xi)$ for $k=0$ or 1 and $\phi_{k}(\xi)=P_{k}(\xi)-P_{k-2}(\xi)$ for $k\geq 2$. Since $\phi_{k}(\pm 1)=0$ for $k\geq 2$, these higher-order modes do not effect the boundary of the element, and hence may be damped by an appropriate filter function. The resulting transform (or Legendre) matrix is given by Eq. (31) in [14].
For a 1D semi-infinite element using LGR points, a similar choice of modal functions is $\phi_{0}(\xi)=e^{-\xi/2}$ and $\phi_{k}(\xi)e^{-\xi/2}\left[L_{k}(\xi)-L_{k-1}(\xi)\right]$. Since $L_{k}(0)=1$ for all $k\geq 0$, we have $\phi_{k}(0)=0$ for $k\geq 1$. The Legendre matrix is then constructed by evaluating $\phi_{k}(\xi_{j})$, where $\xi_{j}$ are the LGR points. Since we employ a tensor product of LGL and LGR points within the semi-infinite element, we construct a corresponding 2D tensor product of modal functions in order to transform the nodal representation. A Boyd-Vandeven filter [5] is applied to the modal representation, and then the solution is inverse transformed to nodal space.
@@ 4 Numerical Results
We assess the efficacy and efficiency of the semi-infinite element approach with a series of one and two-dimensional tests. In 1D, we conduct two tests. The first test evaluates the method’s ability to damp outgoing waves using the wave equation, while the second test is a wave train for the linearized shallow water equations that demonstrates how the absorbing layer handles a series of incoming waves while maintaining the the finite domain solution quality. In 2D, we initially test our method on the advection-diffusion and Helmholtz equations, demonstrating its efficiency and accuracy. For atmospheric simulations, we employ a standard idealized test: the mountain-triggered hydrostatic gravity waves. This test allows us to verify that our approach is effective as an absorbing layer for the compressible Euler equations and showcases its potential for use in atmospheric simulations.
@@
We construct a spectral element method to solve the 1D wave equation as described by (1) for the unknown and flux vectors (2). The space considered is the finite domain $\Omega^{F}=[-2.5,2.5]~{}$m which is discretized using 50 spectral elements of order 6 leading to an effective resolution $\Delta x=0.016~{}$m. One semi-infinite element is added on each end of the finite domain such that the left region is $\Omega^{S}_{L}=(-\infty,-2.5]~{}$m, whereas the right is $\Omega^{S}_{R}=[2.5,\infty)~{}$m. We initialize $u$ and $v$ as follows:
@@
where $x_{c}=0$, and $\sigma=0.15$ m. While the Laguerre functions (SLFs) present an excellent tool for accurately approximating exponentially decaying functions, they are not capable of acting as an absorbing layer on their own. To this end, a sponge, is added to each Laguerre semi-infinite element to damp incoming waves while allowing the SLF basis to accurately approximate this damping. This allows a single semi-infinite element with a sponge to act as the absorbing layer on each side of the domain. We define $X_{L}$ and $X_{R}$, to be the leftmost and rightmost nodes of the domain respectively, such that $X_{R}=11.63$ m, and $X_{L}=-11.63$ m, for semi infinite elements of order 50 and scaling factor of 0.05 m. This test is run for 9 seconds of simulated time with a time step $\Delta t=0.001~{}$s. The Rayleigh damping on each side is controlled by the following profile $\gamma(x)$, adapted from [2], defined for all $x\in\Omega^{S}_{L}\cup\Omega^{S}_{R}$ as follows:
@@
where, ${\rm sgn}$ is the sign function, $X_{0}$ is the coordinate of the first node of the semi-infinite element (in this case $X_{0}=-2.5$ m for the left element and $X_{0}=2.5$ m for the right element). $\Delta\gamma=2\mbox{s}^{-1}$, $L_{0}$ is the end point of each element, such that $L_{0}=X_{L}$ for the left element and $L_{0}=X_{R}$ for the right element, $\alpha=0.3$, and $\zeta=\max(|X_{L}|,|X_{R}|)/18$.
@@
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
@@
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
@@
We define the finite domain as $\Omega^{F}=[0,5000]~{}$m and initialize the fluid to be at rest such that $h(0,x)=0~{}$m and $u(0,x)$. The reference height and velocity are, respectively, $H=10~{}$m and $U=0~{}{\rm ms^{-1}}$. The flow is forced through a time dependent Dirichlet boundary condition at $x=0$ such that $u(t,0)=A\sin(2\pi kt/T)$, where $T=5,000~{}$s is the final time of the simulation, $A=0.025~{}{\rm ms^{-1}}$ is the amplitude of the perturbation, and $k=30$. The finite domain is discretized using 300 spectral elements of order 4, yielding an average spatial resolution of $\Delta x=4.16~{}$m. A Laguerre semi-infinite element is defined as $\Omega^{S}=[5000,\infty)~{}$m and the same damping coefficient defined in (38) is used $\forall x\in\Omega^{S}$, such that $L_{0}=X_{end}$, and $\zeta=L_{0}/18$. The finite domain is discretized using 300 spectral elements of order four, and a Laguerre semi-infinite element of order 50 is used in the damping layer which, with a scaling factor of 100, yields the end point $X_{end}=23,262~{}$m.
As demonstrated in figure 5, the forcing at the boundary instigates a series of wave to move towards the end of the finite domain. Once inside the semi-infinite element, the waves are gradually damped before disappearing entirely. Observation of the finite domain shows that no noticeable changes can be seen in the wave train. To verify this, we perform a simulation of the wave train up to a final time $T_{F}=50,000~{}$s. Figure 6 shows an overlap of the solution taken at an interval $\Delta T=5,000~{}$s starting at $t=0~{}$s; This figure was generated using a low resolution simulation with only $N_{x}=50$ elements in the finite domain to allow the reader to better distinguish the individual points which would be, otherwise, completely overlapping and indistinguishable. The boundary condition at these times is identical and the solution should also be identical as it is periodic in time. It is clear from the figure that there is a remarkable overlap of the finite domain solutions at these times; we can thus conclude that the outgoing waves are effectively damped.
@@
Refer to caption Refer to caption
Refer to caption Refer to caption
@@
Equations (1) with flux and source vectors (5) describe the two-dimensional transport and diffusion of a passive tracer. We performed the transport-diffusion test described in Sec. 4.2.1 of [44] to verify that the coupling between the finite and semi-infinite domain for two dimensional cases. The finite domain extends like $\Omega^{F}=[-5,5]\times[0,10]$, which is subdivided into $N_{x}=12$ elements in the x direction and $N_{z}=125$ elements in the z direction. Both directions use elements of polynomial order 4, for an effective resolution $\{\Delta x,\Delta z\}=\{0.208,0.02\}$ m. The semi-infinite domain, defined as $\Omega^{S}=[-5,5]\times[10,\infty)$, is discretized using 12 Laguerre semi-infinite elements of order 40. Taking into the account a scaling factor $f=0.07$, the last LGR node on each semi-infinite element has a coordinate $Z_{end}=20.09$. To verify that the solution is not adversely affected by inter-domain interface, no damping is used. We initialize the tracer field as follows:
@@
Figure 8 reports the $L_{2}$ and $L_{\infty}$ absolute errors with respect to the exact solution as a function of the $N_{LGL}$ and $N_{LGR}$. We maintain the same number of elements as reported above and vary the order of the elements used in $\Omega^{F}$ and $\Omega^{S}$. Both panels of the figure show that the error floor is determined by the order of the semi-infinite element in the $z$ direction. Increasing the order of elements in $\Omega^{F}$ and in the $x$ direction of $\Omega^{S}$ decreases the error towards this floor. The error values we obtain are also comparable to those reported in table D1 of [44] and this gives us confidence the accuracy of our approach.
@@
We next constructed a spectral element method to solve the governing equation (1) defined by vectors (4) on the 2D semi-infinite channel $\Omega=[0,\infty)\times[-\pi/2,\pi/2]$ subject to a zero Dirichlet BC on the boundary. This case tests the convergence rate of our method and presents some guidance on selecting the orders of the elements in the finite and semi-infinite domains. Applying a nodal SEM using inexact integration to this equation along with the Dirichlet BC yields the matrix-vector problem
@@
$u(x,y)=e^{-x/L}\sin\left(\frac{x}{L}\right)\cos(y)$ (41)
with $L=2$ m and then analytically construct the RHS $f(x,y)$ using the method of manufactured solutions using a wavenumber of $\alpha=10\mbox{m}^{-1}$.
@@
Examining panel b), we see that the choice of the semi-infinite element order $N_{LGR}$ determines a floor for the relative error, while increasing $N_{LGL}$ reduces the relative error to this floor. In particular, choosing $N_{LGR}=48$ and $N_{LGL}=9$ yields a relative error of $\sim 10^{-14}$. Since this problem has a wavelike solution and an exact solution, it provides guidance on choosing the order of the LGL and LGR nodal basis functions. The next test case will show that this approach of using semi-infinite elements can be used to replace the absorbing layer for a benchmark atmospheric flow test case, while yielding good results at a lower cost compared to the standard approach.
@@
We validate our approach on the compressible Euler equations by using it to simulate a classic rising thermal bubble case. In this test a perturbation $\Delta\theta$ is introduced to a neutral atmosphere with constant potential temperature $\theta_{0}=300~{}$K. For our test we consider a finite domain $\Omega^{F}=[-5,5]~{}{\rm km}\times[0,5]~{}{\rm km}$, and discretize it using $N_{z}=N_{x}=20$ elements of order $N=4$ in both directions. This yields a horizontal resolution $\Delta x=125~{}$m and a vertical resolution $\Delta z=62.5~{}$m. The potential temperature perturbation and initial pressure are defined as follows:
$\Delta\theta=\theta_{c}\left(1-\frac{r}{r_{0}}\right)$ (42a)
@@
A set of 20 Laguerre-Legendre semi-infinite elements of order 24 are added at the top of $\Omega^{F}$ yielding the semi-infinite domain $\Omega^{S}=[-5,5]~{}{\rm km}\times[5,\infty)~{}{\rm km}$. To verify that our approach is able to simulate the Euler equations without adversely affecting the solution, we do not use a damping layer for this test. To stabilize the solution past $t=500~{}s$ a viscosity coefficient $\nu=30~{}{\rm m^{2}s^{-1}}$ is used in conjunction with a constant thermal diffusivity $\kappa=2\nu$. This is standard and would be required even for simulations without a semi-infinite element. Free-slip type boundary conditions are used at all domain edges.
The initial potential temperature perturbation generates positive buoyancy, which causes the bubble to rise. Figure 10 shows the solution at $t=1000~{}$s after a significant portion of it has transported through the interface between $\Omega^{F}$ and $\Omega^{S}$ at $z=5$ km. The figure shows that the solution is correctly transported and diffused through the interface. The symmetry of the bubble in the $x$ direction is maintained, and the interface does not introduce any discontinuities. The figure also highlights the improvement in solution quality as the order of the semi-infinite element is increased. The sharpest solution is obtained using a semi-infinite element of order 48 which is displayed in the right panel of figure 10. With our approach validated for the Euler equations, we move on to demonstrating that we can use it for effective absorbing layers in atmospheric flows.
@@
Gravity waves in a stably stratified atmosphere
In this test, a constant horizontal flow impinges on a mountain in a stratified atmosphere. These flow conditions and the mountain’s size determines the structure of the resulting waves. As soon as the flow encounters the mountain, gravity waves quickly propagate both horizontally and vertically. In a proper implementation, these waves should leave the domain with no reflection.
@@
$p=p_{0}\left[1+\frac{g^{2}}{c_{p}\theta_{0}N^{2}}\left(\exp\left(\frac{-zN^{2}}{g}\right)-1\right)\right]^{c_{p}/R}$ (43a)
$\theta=\theta_{0}\exp\left(\frac{zN^{2}}{g}\right),$ (43b)
where $p_{0}=1000~{}$hPa and $\theta_{0}=250~{}$K are the surface values of pressure and potential temperature, and
$N=\frac{g}{\sqrt{c_{p}\theta_{0}}}=0.0196~{}{\rm s^{-1}}$
is the Brunt-Väisälä frequency. An Agnesi mountain with height $h=1~{}$m and half-width $a=10,000~{}$m is located at the center of the domain $x_{c}=0~{}$m with height
$z=\frac{ha^{2}}{(x-x_{c})^{2}+a^{2}}.$
We consider a finite domain $\Omega^{F}=[-120,120]~{}{\rm km}\times[0,15]~{}$km, and discretize it using $N_{x}=120$ elements in the x direction and $N_{z}=21$ elements in the z direction. These elements are of polynomial order 4 in both directions and thus yield the resolutions $(\Delta x,\Delta z)=(500~{}{\rm m},178~{}{\rm m})$. A set of 120 Laguerre-Legendre semi-infinite elements are added at the top of $\Omega^{F}$ yielding the semi-infinite domain $\Omega^{S}=[-120,120]~{}{\rm km}\times[15,\infty)~{}$km. In the horizontal direction, the semi-infinite element uses an order 4 spectral element discretization using LGL nodes, and in the vertical direction it uses an order 14 Laguerre function basis on LGR nodes this yields and average resolution of $\Delta z_{\rm Laguerre}=923~{}$m within $\Omega^{S}$. Due to prevalence of both acoustic waves and gravity waves, as well as the sensitivity of our high order numerical methods, we make use of the spectral low pass Boyd-Vandeven filter [5] to help insure the stability of our simulation (this approach is standard for spectral element simulations of the atmosphere). The filter on the elements in $\Omega^{S}$ is different from that in $\Omega^{F}$ and the continuity of the solution at shared interface nodes is insured by the DSS operation.
To trigger the hydrostatic waves, a uniform eastward wind with speed $U=20.0~{}{\rm ms^{-1}}$ impinges the mountain. We use a sine squared function to define the damping coefficient in the absorbing layer:

### v3

@@ 3.3 Constructing element matrices
$\psi_{l}(\mathbf{x})=h_{i}[\xi(\mathbf{x})]\otimes\overline{h}_{j}[\eta(\mathbf{x})],\hskip 5.0pt\hskip 3.0ptl=i+(j-1)N_{LGL},$ (30)
@@
$\mathbf{M}^{e}_{ij}$ $=\sum_{k=1}^{N_{\xi}}\sum_{m=1}^{N_{\eta}}\overline{\omega}(\xi_{k},\eta_{m})\psi_{i}(\xi_{k},\eta_{m})\psi_{j}(\xi_{k},\eta_{m})|\mathbf{J}(\xi_{k},\eta_{m})|,$ (32)
@@
$\omega_{\xi},\omega_{\eta}$ =$=\begin{cases}\omega,\omega&\textrm{ if $\Omega_{e}\in\Omega^{F}$}\\ \omega,\hat{\omega}&\textrm{ if $\Omega_{e}\in\Omega^{S}$}\end{cases},$ (34)
$N_{\xi},N_{\eta}$ =$=\begin{cases}N_{LGL},N_{LGL}&\textrm{ if $\Omega_{e}\in\Omega^{F}$}\\ N_{LGL},N_{LGR}&\textrm{ if $\Omega_{e}\in\Omega^{S}$}\\ \end{cases}.$ (35)
@@ 3.4 Direct stiffness summation
To control unresolved grid-scale noise and aliasing, low-pass spectral filters are typically used with the spectral element method (SEM) [Sec. 18.3][16]. Spectral filters employ a three-step process to damp/remove unphysical high-frequency components: 1) the element-local nodal solution is transformed into modal space, 2) a low-pass filter is applied in modal space, 3) the filtered modal representation is inverse transformed to nodal space. The nodal to modal transform requires defining a set of modal basis functions. As discussed in Section 3d of [19], the low-pass filter should not violate the continuity requirement of SEM. One way to enforce this is by choosing modal functions such that most of the modes are zero at the element boundary. For a 1D element using LGL points, an appropriate choice is $\phi_{k}(\xi)=P_{k}(\xi)$ for $k=0$ or 1 and $\phi_{k}(\xi)=P_{k}(\xi)-P_{k-2}(\xi)$ for $k\geq 2$. Since $\phi_{k}(\pm 1)=0$ for $k\geq 2$, these higher-order modes do not effect the boundary of the element, and hence may be damped by an appropriate filter function. The resulting transform (or Legendre) matrix is given by Eq. (31) in [19].
For a 1D semi-infinite element using LGR points, a similar choice of modal functions is $\phi_{0}(\xi)=e^{-\xi/2}$ and $\phi_{k}(\xi)e^{-\xi/2}\left[L_{k}(\xi)-L_{k-1}(\xi)\right]$. Since $L_{k}(0)=1$ for all $k\geq 0$, we have $\phi_{k}(0)=0$ for $k\geq 1$. The Legendre matrix is then constructed by evaluating $\phi_{k}(\xi_{j})$, where $\xi_{j}$ are the LGR points. Since we employ a tensor product of LGL and LGR points within semi-infinite element, we construct a corresponding 2D tensor product of modal functions in order to transform the nodal representation. A Boyd-Vandeven filter [5] is applied to the modal representation, and then the solution is inverse transformed to nodal space.
@@ 4 Numerical Results
We assess the efficacy and efficiency of the semi-infinite element approach with a series of one and two-dimensional tests. In 1D, we conduct two tests. The first test evaluates the method’s ability to damp outgoing waves using the wave equation, while the second test is a wave train for the linearized shallow water equations that demonstrates how the absorbing layer handles a series of incoming waves while maintaining the finite domain solution quality. In 2D, we initially test the proposed method on the advection-diffusion and Helmholtz equations. For atmospheric simulations, we employ three standard idealized tests: the mountain-triggered hydrostatic gravity waves over a single mountain, and the non-hydrostatic wave problem triggered by a Scär ridge. These tests –with increasing physical complexity– allow us to verify that the proposed approach is effective as an absorbing layer for the compressible Euler equations and showcase its potential for use in atmospheric simulations.
@@
We solve the 1D wave equation as described by (1) for the unknown and flux vectors (3). The space considered is the finite domain $\Omega^{F}=[-2.5,2.5]~{}$m which is discretized using 50 spectral elements of order 6 leading to an effective resolution $\Delta x=0.016~{}$m. One semi-infinite element is added on each end of the finite domain such that the left region is $\Omega^{S}_{L}=(-\infty,-2.5]~{}$m, whereas the right is $\Omega^{S}_{R}=[2.5,\infty)~{}$m. We initialize $u$ and $v$ as
@@
where $x_{c}=0$ and $\sigma=0.15$ m. While the scaled Laguerre functions (SLFs) present an excellent tool for accurately approximating exponentially decaying functions, they are not capable of acting as an absorbing layer on their own. To this end, a sponge is added to each Laguerre semi-infinite element to damp incoming waves while allowing the SLF basis to accurately approximate this damping. This allows a single semi-infinite element with a sponge to act as the absorbing layer on each side of the domain. We define $X_{L}$ and $X_{R}$ to be the leftmost and rightmost nodes of the domain respectively, such that $X_{L}=-11.63$ m and $X_{R}=11.63$ m for semi infinite elements of order 50 and a scaling factor $\lambda=0.05~{}$m. This test is run for 9 seconds of simulated time with a time step $\Delta t=0.001~{}$s. The Rayleigh damping on each side is controlled by $\gamma(x)$, adapted from [2], defined for all $x\in\Omega^{S}_{L}\cup\Omega^{S}_{R}$ as follows:
@@
where ${\rm sgn}$ is the sign function, $X_{0}$ is the coordinate of the first node of the semi-infinite element (in this case $X_{0}=-2.5$ m for the left element and $X_{0}=2.5$ m for the right element). $\Delta\gamma=2\mbox{~{}s}^{-1}$, $L_{0}$ is the end point of each element, such that $L_{0}=X_{L}$ for the left element and $L_{0}=X_{R}$ for the right element, $\alpha=0.3$, and $\zeta=\max(|X_{L}|,|X_{R}|)/18$.
@@
We define the finite domain as $\Omega^{F}=[0,5000]~{}$m and initialize the fluid to be at rest such that $h(0,x)=0~{}$m and $u(0,x)=0~{}{\rm ms^{-1}}$. The reference height and velocity are, respectively, $H=10~{}$m and $U=0~{}{\rm ms^{-1}}$. The flow is forced through a time dependent Dirichlet boundary condition at $x=0$ such that $u(t,0)=A\sin(2\pi kt/T)$, where $T=5,000~{}$s is the final time of the simulation, $A=0.025~{}{\rm ms^{-1}}$ is the amplitude of the perturbation, and $k=30$. The finite domain is discretized using 300 spectral elements of order 4, yielding an average spatial resolution of $\Delta x=4.16~{}$m. A Laguerre semi-infinite element is defined as $\Omega^{S}=[5000,\infty)~{}$m and the same damping coefficient defined in (40) is used $\forall x\in\Omega^{S}$, such that $L_{0}=X_{end}$, and $\zeta=L_{0}/18$. A Laguerre semi-infinite element of order 50 is used in the damping layer which, with a scaling factor $\lambda=100~{}$m, results in the end point $X_{end}=23,262~{}$m.
As demonstrated in figure 5, the forcing at the boundary instigates a series of waves to move towards the end of the finite domain. Once inside the semi-infinite element, the waves are gradually damped before disappearing entirely. Observation of the finite domain shows that no noticeable changes can be seen in the wave train. To verify this, we perform a simulation of the wave train up to a final time $T_{F}=50,000~{}$s. Figure 6 shows an overlap of the solution taken at an interval $\Delta T=5,000~{}$s starting at $t=0~{}$s; This figure was generated using a low resolution simulation with only $N_{x}=50$ elements in the finite domain to allow the reader to better distinguish the individual points which would be, otherwise, completely overlapping and indistinguishable. The boundary condition at these times is identical and the solution should also be identical as it is periodic in time. It is clear from the figure that there is a remarkable overlap of the finite domain solutions at these times; we can thus conclude that the outgoing waves are effectively damped.
@@
Equations (1) with flux and source vectors (6) describe the two-dimensional transport and diffusion of a passive tracer. We performed the transport-diffusion test described in Sec. 4.2.1 of [52] to verify the coupling between the finite and semi-infinite domain for two dimensional cases. The finite domain extends like $\Omega^{F}=[-5,5]\times[0,10]$, which is subdivided into $N_{x}=12$ elements in the x direction and $N_{z}=125$ elements in the z direction. Both directions use elements of polynomial order 4, for an effective resolution $\{\Delta x,\Delta z\}=\{0.208,0.02\}$ m. The semi-infinite domain, defined as $\Omega^{S}=[-5,5]\times[10,\infty)$, is discretized using 12 Laguerre semi-infinite elements of order 40. Taking into account a scaling factor $\lambda=0.07~{}$m, the last LGR node on each semi-infinite element has a coordinate $Z_{end}=20.09$. To verify that the solution is not adversely affected by inter-domain interface, no damping is used for this test. We initialize the tracer field as follows:
@@
Figure 8 reports the $L_{2}$ and $L_{\infty}$ absolute errors with respect to the exact solution as a function of the $N_{LGL}$ and $N_{LGR}$. We maintain the same number of elements as reported above and vary the order of the elements used in $\Omega^{F}$ and $\Omega^{S}$. Both panels of the figure show that the error floor is determined by the order of the semi-infinite element in the $z$ direction. Increasing the order of elements in $\Omega^{F}$ and in the $x$ direction of $\Omega^{S}$ decreases the error towards this floor. The error values we obtain are also comparable to those reported in table D1 of [52] and this gives us confidence the accuracy of this approach.
@@
We Helmholtz equation is obtained from (1) by means of vectors (5). We solved it on the 2D semi-infinite channel $\Omega=[0,\infty)\times[-\pi/2,\pi/2]$ subject to a zero Dirichlet boundary conditions (BCs) on the boundary. This test verifies the convergence rate of the solution and presents some guidance on selecting the orders of the elements in the finite and semi-infinite domains. Applying nodal SEM with inexact integration to this equation along with the Dirichlet BCs yields the matrix-vector problem
@@
$u(x,y)=e^{-x/L}\sin\left(\frac{x}{L}\right)\cos(y),$ (43)
with $L=2$ m and then analytically construct the RHS $f(x,y)$ using the method of manufactured solutions using a wavenumber $\alpha=10\mbox{m}^{-1}$.
@@
Close examination of panel b) shows that the choice of the semi-infinite element order $N_{LGR}$ determines a floor for the relative error, while the increase of $N_{LGL}$ reduces the relative error to this floor. In particular, choosing $N_{LGR}=48$ and $N_{LGL}=9$ yields a relative error of $\sim 10^{-14}$. Since this problem has a wavelike solution and an exact solution, it provides guidance on choosing the order of the LGL and LGR nodal basis functions. The next test case will show that semi-infinite elements can be used in the absorbing layer for reducing the cost in a benchmark atmospheric flow test case.
@@
We validate the Laguerre-Legendre semi-infinite element approach on the compressible Euler equations by using it to simulate a classic rising thermal bubble case. In this test a perturbation $\Delta\theta$ is introduced to a neutral atmosphere with uniform potential temperature $\theta_{0}=300~{}$K. The finite domain is $\Omega^{F}=[-5,5]~{}{\rm km}\times[0,5]~{}{\rm km}$, which is subdivided into $N_{z}=N_{x}=20$ elements of order $N=4$. This yields a horizontal resolution $\Delta x=125~{}$m and a vertical resolution $\Delta z=62.5~{}$m. The potential temperature perturbation and initial pressure are defined as follows:
$\Delta\theta=\theta_{c}\left(1-\frac{r}{r_{0}}\right)\mbox{ if }|r|\leq r_{0}$ (44a)
@@
A set of 20 Laguerre-Legendre semi-infinite elements of order 24 is added on top of $\Omega^{F}$ to build the semi-infinite domain $\Omega^{S}=[-5,5]~{}{\rm km}\times[5,\infty)~{}{\rm km}$. To verify that this approach is able to solve the Euler equations without adversely affecting the solution, we do not use a damping layer for this test and simply verify that the rising thermal crosses the $\Omega^{F}-\Omega^{S}$ boundary without being affected by the change in element type and drastic change in resolution. To stabilize the solution past $t=500~{}s$ a viscosity coefficient $\nu=30~{}{\rm m^{2}s^{-1}}$ is used in conjunction with a constant thermal diffusivity $\kappa=2\nu$. This is standard and would be required even for simulations without a semi-infinite element. Free-slip type boundary conditions are used at all domain edges.
The initial potential temperature perturbation generates positive buoyancy, which causes the bubble to rise. Figure 10 shows the solution at $t=1000~{}$s after a significant portion of it has transported through the interface between $\Omega^{F}$ and $\Omega^{S}$ at $z=5$ km. The figure shows that the solution is correctly transported and diffused through the interface. The symmetry of the bubble in the $x$ direction is maintained, and the interface does not introduce spurious noise or discontinuities. The figure also highlights the improvement in solution quality as the order of the semi-infinite element is increased. The sharpest solution is obtained using a semi-infinite element of order 48 which is displayed in the right panel of figure 10. With this approach validated for the Euler equations, we move on to demonstrating that we can use it for effective absorbing layers in atmospheric flows.
@@
Linear hydrostatic mountain waves
In this test, a constant horizontal flow with velocity $U=20~{}\rm{ms^{-1}}$ impinges on a mountain in a stratified atmosphere. These flow conditions and the mountain’s size determines the structure of the resulting waves. As soon as the flow encounters the mountain, gravity waves quickly propagate both horizontally and vertically. In a proper implementation, these waves should leave the domain with no reflection.
@@
$p=p_{0}\left[1+\frac{g^{2}}{c_{p}\theta_{0}N^{2}}\left(\exp\left(\frac{-z\mathcal{N}^{2}}{g}\right)-1\right)\right]^{c_{p}/R}$ (45a)
$\theta=\theta_{0}\exp\left(\frac{z\mathcal{N}^{2}}{g}\right),$ (45b)
where $p_{0}=1000~{}$hPa and $\theta_{0}=250~{}$K are the sea level values of pressure and potential temperature, and
$\mathcal{N}=\frac{g}{\sqrt{c_{p}\theta_{0}}}=0.0196~{}{\rm s^{-1}}$
is the Brunt-Väisälä frequency. An Agnesi mountain with height $h=1~{}$m and half-width $a=10,000~{}$m is located at the center of the domain $x_{c}=0~{}$m with shape
$z=\frac{ha^{2}}{(x-x_{c})^{2}+a^{2}}.$ (46)
For this test as well as for the other case involving topography, a terrain following sigma coordinate [14] is used. We consider a finite domain $\Omega^{F}=[-120,120]~{}{\rm km}\times[0,15]~{}$km, and subdivided it into $N_{x}\times N_{z}=120\times 21$ elements of order 4, which lead to the resolution $(\Delta x,\Delta z)=(500~{}{\rm m},178~{}{\rm m})$. A set of 120 Laguerre-Legendre semi-infinite elements are added on top of $\Omega^{F}$ yielding the semi-infinite domain $\Omega^{S}=[-120,120]~{}{\rm km}\times[15,\infty)~{}$km. In the horizontal direction, each semi-infinite element uses an order 4 spectral element discretization using LGL nodes, and in the vertical direction it uses an order 14 Laguerre function basis on LGR nodes and a scaling factor $\lambda=300~{}$m which yields an end point $Z_{end}=28853$. Due to prevalence of both acoustic waves and gravity waves, as well as the sensitivity of high order numerical methods, we make use of the spectral low pass Boyd-Vandeven filter [5] to help insure the stability of the simulation (this approach is standard for spectral element simulations of the atmosphere). The filter on the elements in $\Omega^{S}$ is different from that in $\Omega^{F}$ and the continuity of the solution at shared interface nodes is insured by the DSS operation.
We use a sine squared function to define the damping coefficient in the vertical absorbing layer: